OpenStudy (anonymous):

The width of a rectangle is two-thirds of its length, and its area is 216 square meters. What is the length? Step by step

OpenStudy (anonymous):

Let length = x then Width = \[(2/3)x\] Area = lw = \[(2/3)x^2\] \[Area = lw = (2/3)x^2= 216 cm^2\] Can you go on from there?

OpenStudy (anonymous):

Sorry, units are m not cm